OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

Directrix (directrix):

The slope of the line will be -3 because slopes of parallel lines are equal. Use the point-slope form of a line to get the equation. y - y1 = m(x - x1) where m = -3 and (x1,y1) = (-1,6) y - 6 = -3(x - (-1)) y - 6 = -3(x + 1) y - 6 = -3x - 3 y = -3x - 3 + 6 y = -3x + 3